### v3

@@ Appendix I Heterogeneous treatment effects
Treatment ${}^{***}$ (0.265) ${}^{***}$ (0.239)
Social science --0.124 (0.125)
STEM ${}^{***}$ (0.099) ${}^{***}$ (0.111)
Family friends IT ${}^{***}$ (0.092)
Small Town $*{}^{*}$ (0.102) $*{}^{*}$ (0.126)
Warsaw 0.088 (0.094) ${}^{***}$ (0.137)
Mother ${}^{***}$ (0.137)
Applied before ${}^{***}$ (0.096)
First time mentor 0.123 (0.084)
Managerial experience ${}^{***}$ (0.087)
Career changer ${}^{***}$ (0.099)
Years of experience mentor --0.003 (0.013)
Non-traditional credential ${}^{***}$ (0.188) ${}^{***}$ (0.179)
Masters ${}^{***}$ (0.171) 0.218 (0.135)
No education mentioned ${}^{**}$ (0.193) $*{}^{*}$ (0.287)
Postgraduate $*{}^{*}$ (0.188) ${}^{***}$ (0.194)
Age: 25-30 $*{}^{*}$ (0.145) ${}^{**}$ (0.155)
Age 31-35 ${}^{***}$ (0.168) 0.170 (0.171)
Age: 36+ ${}^{**}$ (0.198) --0.144 (0.184)
Prof. experience ${}^{***}$ (0.013) 0.008 (0.011)
Treatment×$\times$×Social science 0.273 (0.179)
Treatment×$\times$×STEM ${}^{***}$ (0.136) ${}^{***}$ (0.160)
Treatment×$\times$×Family friends IT $*{}^{*}$ (0.119)
Treatment×$\times$×Small ${}^{***}$ (0.131) ${}^{***}$ (0.173)
Treatment×$\times$×Warsaw ${}^{***}$ (0.129) ${}^{***}$ (0.182)
Treatment×$\times$×Mother ${}^{***}$ (0.164)
Treatment×$\times$×Applied before --0.155 (0.123)
Treatment×$\times$×First time mentor --0.079 (0.108)
Treatment×$\times$×Managerial experience ${}^{***}$ (0.112)
Treatment×$\times$×Career changer ${}^{***}$ (0.125)
Treatment:$\times$×Years of experience mentor 0.001 (0.016)
Treatment×$\times$ Non-traditional credential ${}^{***}$ (0.235) ${}^{***}$ (0.251)
Treatment×$\times$ masters ${}^{***}$ (0.209) --0.267 (0.192)
Treatment×$\times$ No education mentioned ${}^{***}$ (0.347) --0.139 (0.349)
Treatment×$\times$×postgraduate ${}^{**}$ (0.236) ${}^{***}$ (0.243)
Treatment×$\times$×Age: 25-30 ${}^{***}$ (0.203) ${}^{***}$ (0.209)
Treatment×$\times$×Age:31-35 ${}^{***}$ (0.241) ${}^{***}$ (0.234)
Treatment×$\times$×Age: 36+ 0.273 (0.277) --0.013 (0.235)
Treatment×$\times$×Prof experience ${}^{***}$ (0.016) $*{}^{*}$ (0.015)
Observations 4,440 4,965
[…]
Log Likelihood --11,845.370 --6,329.277
Wald Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
LR Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Score (Logrank) Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
Note: Estimates from the Cox proportional hazard model. Outcome variable tech job.
Appendix J Models used in off-policy sections
This section presents further details about the propensity and outcome models used in Section 6 for policy estimation. To estimate both models, we first join the Mentoring, Challenges, and applicants’ datasets. We obtain a dataset of 1314 individuals who applied to Dare IT, out of whom 151 participated in Mentoring and 166 in Challenges; the remaining 997 are the non-selected Mentoring applicants, late applicants to Challenges, and the control groups. All individuals are observed over the span of 15 months. The dataset is randomly split between training and testing at the individual applicant level, 50-50.
Propensity model.
To estimate the propensity model for policy estimation, we use multinomial logit regression with the full set of covariates that overlap between all datasets. We train the model in the training sample and predict into the testing set. We find that the propensity of being assigned to Mentoring ranges from 4% to 37%, to Challenges from 5% to 32%, and to Out of Dare IT from 53% to 89%. Figure 13 presents the histogram of the three propensities.
Figure 13: Predicted program propensities.
Note: Test set propensity of assignment across groups, in blue to Challenges, in red to Mentoring, and in green to Out of Dare IT.
We also evaluate the adjustment with the propensity weights by compare the standardized absolute mean differences in covaraites values with and without the adjustment. In Figure 14, we show the difference in average values of covariates between Mentoring and Nothing groups, with and without adjustment in each cross-fitting round. The same analysis for the case of Challenges and Nothing is in Figure 15. We find that the propensity model successfully adjusts for the differences in covariates.
Figure 14: Covariate balance: Mentoring.
Note: Differences in mean covariates across Mentoring and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Figure 15: Covariate balance: Challenges.
Note: Differences in mean covariates across Challenges and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Outcome model.
We use Cox proportional hazard model to model outcomes. We estimate a model with a full set of covariates and interactions of covariates and treatments. We model outcomes at the person and month level. In Figure 16, we show an evaluation of the model’s predictive performance. First, we train a model on the training sample and predict probabilities of having a job in a specific month using the testing sample; then, we take the average per user and group users based on these averages into quintiles. Finally, we compare the predicted probability within the quintile to the observed probability. We find that the differences are statistically insignificant, and the magnitudes are small. We conclude that the model is well-calibrated.
Figure 16: Comparison of predicted and observed outcomes.
Note: Comparison of predicted and observed outcomes in the test set. Subjects are grouped into quintiles based on the predicted probability of having a tech job. In red, the average per quintile predicted probability. In blue, is the average per quintile observed outcome. Whiskers show confidence intervals.
J.1 Cross-fitting.
We estimate predicted outcomes under a policy using cross-fitting with 10 folds and 100 repetitions. Specifically, for each repetition, $m\in\left\{1,...,100\right\}$, we randomly assign observations into 10 folds $k\in\left\{1,..,10\right\}$. Randomization is at the individual subject level with equal probability. Next, for each fold $k$ we train an outcome and propensity models using data from all other folds and predict outcomes and treatments in the $k$-th fold: $\hat{\mu}_{k}(X_{i})_{P}^{m}$ and $\hat{e}_{k}(X_{i})_{P}^{m}$, where $P\in\left\{0,M,C\right\}$ refers to a program. We use these predictions to construct AIPW estimates; for example, the treatment effect from Mentoring instead of Out of Dare IT writes:
$\tau^{m}_{OM}=\frac{1}{n}\sum_{i=1}^{n}\left[\hat{\mu}_{k}(X_{i})_{M}^{m}+\frac{\mathbb{1}_{W_{i}=M}}{\hat{e}_{k}(X_{i})^{m}_{M}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{M})-\hat{\mu}(X_{i})_{0}^{m}+\frac{\mathbb{1}_{W_{i}=0}}{\hat{e}_{k}(X_{i})^{m}_{O}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{O})\right],$
analogously for Challenges. Finally, following Chernozhukov et al. (2017), we take the median over 100 repetitions so that $\hat{\tau}_{OM}=Median(\tau^{m}_{0M})$.
Appendix K Policy estimation algorithms
@@ 2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix J.1 for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@
Appendix L Validate AIPW approach
In Section 6, we used estimated AIPW scores to construct and evaluate counterfactual policies. Now, we compare the estimates of treatment effects using the AIPW approach in the testing set with cross-fitting and the experimental results. Thus, we want to test the hypotheses that $\tau^{AIPW}=\tau^{RCT}$. To estimate $\tau^{AIPW}$, we proceed in the same steps as in Section 6: first, we pool all datasets: Mentoring treatment and control, mentoring non-selected applicants, Challenges treatment and control, challenges late applicants. We split the dataset into train and test, and in the testing dataset, we randomly assign 10 folds. Then for each fold, we train a proportional hazard model for the outcomes and a multinomial logistic regression for propensities using other folds, and then we predict outcomes and propensities for the left-out fold and construct AIPW scores. In the outcome model, we include an indicator variable of whether the applicant was selected by a mentor or not. We repeat the process of cross-fitting 100 times and take the average AIPW score per observation.
To assess Mentoring, we take the difference between predicted AIPW scores under mentoring treatment and under mentoring control, i.e., the estimate of the average treatment effect for the entire sample, analogously for Challenges. We compare obtained quantities with the estimates of the treatment effects using Cox models on the experimental samples. We also compute the difference between the estimates. Figure 17 shows the results
Figure 17: Difference between AIPW and RCT estimates.
Note: Comparison of the average treatment effect estimates using the randomized experiment and the AIPW scores. Point estimates and confidence intervals.
We find that the RCT and AIPW estimates are similar, and the differences are statistically not distinguishable from zero. Recall that the randomization into treatment and control was carried out within experimental samples, that is, the sample of Mentoring applicants who mentors selected and the subset of applicants to Challenges that applied to the program before we closed the registration. Thus, it was not a priori clear that the AIPW procedure with cross-fitting would recover similar estimates of the treatment effect after we pooled all datasets together. The finding that these estimates are very close in magnitude and statistically not different from one another bolsters the credibility of results presented in Section 6.
Appendix M Mentors’ selection strategies
In this section, we dive deeper into four mentors’ strategies: Observed, Highest CATE, Highest if treated, and Least promising. We analyze the degree to which they overlap in terms of which candidates get selected and how the characteristics of potential mentees differ across these strategies.
First, we present the Venn diagram in Figure 18. We observe considerable overlap between Observed, Highest CATE, Highest if treated; approximately half of the candidates selected for Mentoring under Highest CATE are also selected in the observed allocation. Furthermore, almost two-thirds of selections under Highest if treated and Highest CATE overlap. Similarly, 100 out of 138 candidates overlap between Highest if trated and Observed selection. In contrast, there is low overlap between Least promising and the other three policies, with 124 candidates that are selected under Least promising not being selected under any other policy.
Figure 18: Venn diagram of selected policies.
Note: Each policy has 138 selected candidates. Percentages refer to the share of the 138 that fall into the region of the graph.
Next, we consider the following sets of candidates: (i) Observed selection – all candidates selected under the current allocation, (ii) Only least promising – candidates that are selected under the Least promising policy and are not selected under the remaining three policies, (iii) Highest CATE & Observed selection – the candidates that are selected both under the Highest CATE and under Observed selection, and (iv) Highest CATE but not selected - where we consider candidates that are selected under Highest CATE, but under mentors’ rule. We analyze the baseline outcomes, average treatment effects, and means of selected covariates. We present results in Table 11
Table 11: ATEs, baselines, and covariate means and standard errors for different policies
policies ATE Baseline STEM Prof. exp 30 plus Small town
Observed selection $0.095(0.012)$ $0.275(0.008)$ $0.457(0.011)$ $7.072(0.110)$ $0.514(0.168)$ $0.232(0.009)$
Only least promising $0.125(0.008)$ $0.048(0.005)$ $0.076(0.006)$ $6.277(0.160)$ $1.000(0.000)$ $0.588(0.012)$
Highest CATE & Observed selection $0.158(0.017)$ $0.244(0.011)$ $0.493(0.015)$ $8.587(0.161)$ $0.507(0.015)$ $0.400(0.015)$
Highest CATE but not selected $0.208(0.016)$ $0.112(0.011)$ $0.365(0.016)$ $10.095(0.168)$ $0.365(0.016)$ $0.873(0.011)$
Note: Average treatment effects and baselines from Section 6.2 standard errors clustered at the candidate level. STEM is the share of candidates selected under the considered policy that have a degree in STEM, Prof. exp is the average length of professional experience, 30 plus is the share of candidates that are above 30 years old, and Small town is the share of candidates that reside in a smaller town.
Comparing candidates selected only under Least promising to other categories, we find that these applicants are less likely to have a STEM degree, have shorter professional experience, and are more likely to be above 30 years old. Even though both mentors overlook this group and we do not select these candidates under the Highest CATE policy, the estimated treatment effects for this group are high. Furthermore, the treatment effect of 0.125 increase in the probability of getting a tech job is particularly high considering the baseline of only 0.048.
Finally, comparing the candidates that have high estimated treatment effects and get selected to those with high treatment effects who are not selected, we notice that first, mentors select high CATE candidates with higher baselines and lower average treatment effects. They tend not to select enough candidates who have longer professional experience and come from smaller towns. These patterns are consistent with our estimates of the heterogenous treatment effects shown in Section 5.2.

### v4

@@ Appendix G Bounds on the Average Treatment Effect Estimates
   ATE 0.153 0.089
   SE (0.057) (0.043)
   ATE % baseline 33.7 45.3
   SE % baseline (12.6) (21.8)
   N Treatment 151 183
   N Control 147 225
Note: Estimates of the average treatment effect of Mentoring and Challenges at 4 months after the end of each program. Estimates obtained using the difference-in-means estimator.
Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition. The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts card2018works. If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
Appendix H Policy estimation algorithms
@@ 2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix LABEL:cross_fit_appendix for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@
H.1 Assignment Models
We pool data from Mentoring experimental participants (N=298), Challenges experimental participants (N=408), non-selected Mentoring applicants (N=468), and late Challenges applicants (N=160), yielding 1,334 total applicants. All individuals are observed over fifteen months post-application. We randomly split this dataset 50-50 at the applicant level into training (N=667) and testing (N=667) sets.
Using the training set, we estimate predicted treatment effects for each program relative to non-participation. We fit two causal survival forests (athey2019generalized; cui2023estimating): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. These settings follow standard recommendations in the generalized random forests (grf) package (tibshirani2024grf).
The causal survival forest estimates conditional average treatment effects on the restricted mean survival time (RMST) scale—the average time until technology employment over our observation horizon. More negative predicted effects indicate faster job finding under treatment. We apply the trained models to the test set to obtain predicted treatment effects $\hat{\tau}_{i}^{M}$ (Mentoring vs. nothing) and $\hat{\tau}_{i}^{C}$ (Challenges vs. nothing) for each test-set applicant.
Appendix I Evaluation Models
While assignment models determine which applicants receive which program under various policies, evaluation models estimate the causal effects of these assignments. The credibility of our policy comparisons rests on these evaluation models accurately estimating treatment effects for the allocated applicants. If evaluation models correctly estimate treatment effects, our policy value estimates remain valid even if assignment models imperfectly rank applicants.
We use the held-out test set (N=667) for all policy evaluations. This sample includes experimental participants and non-participants from both programs, ensuring we can estimate counterfactual outcomes under different assignment policies. All individuals are observed over the same fifteen-month post-application horizon.
Model specification.
For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (athey2019generalized; cui2023estimating). Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
Each causal survival forest uses identical parameters to the assignment models: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. Model fitting follows the grf package implementation (tibshirani2024grf).
The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively. The forest automatically constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a survival forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates. These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
Cross-fitting procedure.
To avoid overfitting and ensure valid inference, we employ repeated $K$-fold cross-fitting on the test set. Specifically, for each repetition $m\in\{1,\ldots,100\}$, we randomly assign observations into $K=5$ folds. Randomization is stratified by experimental group (Mentoring treated, Mentoring control, Challenges treated, Challenges control, non-participants) to ensure each fold contains observations from all groups. For each fold $k\in\{1,\ldots,5\}$, we:
  1. 1.
    Train causal survival forests using data from all folds except $k$ (the training sample for fold $k$).
  2. 2.
    Predict AIPW scores for observations in fold $k$: $\psi_{i,k}^{M,m}$ and $\psi_{i,k}^{C,m}$ for each individual $i$ in fold $k$, along with corresponding variance estimates $\hat{V}_{i,k}^{M,m}$ and $\hat{V}_{i,k}^{C,m}$.
After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual. We then aggregate across the 100 repetitions: following chernozhukov2017double, we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$. Variance estimates are similarly aggregated by taking the median across repetitions.
Policy value estimation.
Given an allocation policy $\pi$ that assigns each test-set individual to Mentoring, Challenges, or neither, we estimate the policy value as:
$\hat{V}(\pi)=\frac{1}{n_{M}}\sum_{i:\pi(i)=M}\psi_{i}^{M}+\frac{1}{n_{C}}\sum_{i:\pi(i)=C}\psi_{i}^{C}$
where $n_{M}$ and $n_{C}$ are the number of individuals assigned to Mentoring and Challenges, respectively. Standard errors are computed using the aggregated variance estimates from the causal survival forests, accounting for the within-arm averaging. For comparisons between policies or programs, we compute appropriate differences in AIPW scores and propagate variances assuming independence across programs.
This approach separates the assignment decision (which applicants receive which program) from the evaluation of that decision (estimating treatment effects for the assigned groups). The validity of our policy comparisons depends on the evaluation models accurately estimating treatment effects, not on the assignment models perfectly ranking applicants.
I.1 Model Validation: Recovering Experimental Effects
To validate our evaluation model specification, we assess whether cross-fitted AIPW scores correctly recover experimental average treatment effects. This diagnostic tests whether our models produce unbiased treatment effect estimates in the setting where we have the benchmark from the randomized experiment.
Validation procedure.
We train causal survival forests following the same process as before, using 5-fold cross-fitting and generate out-of-sample AIPW predictions for each participant; we use the experimental samples. We compare these individual-level predictions to the experimental ATE by examining the distribution of deviations: $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$. Under correct model specification, these deviations should center at zero with dispersion reflecting genuine treatment effect heterogeneity plus estimation noise.
Figure 14 presents the results. The red solid vertical line indicates the RCT benchmark (zero deviation), with shaded regions marking 95% confidence intervals around both the RCT estimate (red) and the mean AIPW deviation (gray). The black dashed line shows the mean deviation from the RCT estimate.
Figure 14: Model Validation: Distribution of Deviations from RCT Estimates
Refer to caption Refer to caption
(a) Mentoring Program (b) Challenges Program
Notes: Distribution of deviations between cross-fitted AIPW predictions and experimental ATE benchmarks. Each panel shows $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$ for experimental participants. The red vertical line marks zero deviation (perfect recovery), with red shading indicating the 95% confidence interval around the RCT estimate. The black dashed line shows the mean AIPW deviation, with gray shading for its 95% confidence interval. Causal survival forests trained exclusively on experimental samples using 5-fold cross-fitting. Mentoring: mean deviation = $0.08$ months (SE = 0.03); Challenges: mean deviation = $0.06$ months (SE = 0.04). Deviations statistically indistinguishable from zero in both programs.
For the Mentoring program (Panel a), AIPW predictions center tightly around the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03). The distribution is approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE. The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
For the Challenges program (Panel b), AIPW predictions show similarly strong performance, with a mean deviation of $0.06$ months (SE = 0.04). While the distribution exhibits slightly more dispersion than Mentoring—consistent with Challenges’ larger experimental sample and potentially greater heterogeneity—the mean deviation remains statistically indistinguishable from zero. The slight positive mean suggests modest underestimation, but the bias is economically negligible (less than 2% of the experimental ATE).
The cross-fitted causal survival forests accurately recover experimental treatment effects, demonstrating that our evaluation approach is well-calibrated and unbiased within the experimental sample. This gives us confidence that the models correctly estimate treatment effects where treatment is randomized.